# Robust Robotic Exploration and Mapping Using Generative Occupancy Map Synthesis

### v1

@@ 4 SceneSense Design and Implementation
SceneSense [13, 14] is a diffusion model designed to generate 3D occupancy predictions given an observed occupancy map. Notably, SceneSense only generates occupancy predictions in space that has not been observed, and therefore maintains the fidelity of the observed map. In this section we will discuss the design of the SceneSense network as an unconditional diffusion model as well as our method for training. Then we will discuss our inference time control method of occupancy inpainting.
@@ 4.1.1 Denoising Network
We design our denoising network as an unconditional diffusion model. Specifically the denoising network is a U-net constructed from the huggingface diffusers library of blocks [41] and consists of Resnet [42] downsampling/upsampling blocks. Notably, as discussed in Reed et al. [13, 14] the addition of conditioning for guidance as often seen in text-to-image diffusion models [10] does not generate more accurate predictions and increases inference times in this particular application. These ablations are presented in Appendix A and B.
@@ 4.1.2 Occupancy Mapping
Occupancy mapping allows for platforms to build a running map of areas that have been measured to contain matter using onboard sensors like lidar or RGB-D cameras. For our framework we use the popular occupancy mapping framework OctoMap [17] to generate an occupancy map as the platform explores the environment. Importantly, OctoMap provides a probability of occupancy $o\in[0,1]$ for every voxel in the map that has been observed using pose ray casting. This means that as we explore we will be maintaining not only a map of occupied areas $M_{o}$, but also a map of areas that have been measured to not contain any data $M_{u}$. These maps will later be used to inform SceneSense where occupancy predictions should be made.
@@ 4.2.1 Data Augmentation
[…]
@@ 4.2.2 Occupancy Corruption
To corrupt each ground truth local occupancy map $x$ to train the network we add Gaussian noise to $x$ to generate $x_{t}$. This corruption process is defined in 3 where the intensity of the noise is controlled by $\alpha_{t}$ which is configured by a linear noise scheduler [31].
@@ 4.2.3 Loss Function
The network $f_{\theta}$ is trained using the calculated $l_{2}$ loss between the denoised $x_{t}$ prediction and the associated ground truth data $x$. $l_{2}$ loss is a popular diffusion loss function, however other loss functions such as cross-entropy loss or mean squared error can be applied and have had some success in similar diffusion frameworks [43, 10, 44, 45].
@@ 4.3.2 Occupancy Inpainting
Our method of occupancy inpainting is provided in Algorithm 1 which ensures observed space is never overwritten with SceneSense predictions. Inspired by image inpainting methods seen in image diffusion [10] and guided image synthesis methods [46], occupancy inpainting continuously applies the known occupancy information to the diffusion target during inference. To perform occupancy inpainting we select our target region $x$ as a submap of the full occupancy map. From $x$ we build an occupied map $M_{o}$ and an unoccupied map $M_{u}$ composed of observed occupied voxels and observed unoccupied voxels, respectively. From $M_{u}$ and $M_{o}$ we generate noisy representations commensurate with the current diffusion step $t$ using Eq. 3, $M_{un}$ and $M_{on}$, respectively. Finally the data from $M_{un}$ and $M_{on}$ replaces the data in the diffusion target at the associated coordinates.
@@
5:      $x_{obs}^{t-1}\sim\mathcal{N}\left(\sqrt{\alpha_{t-1}}x_{obs},(1-\alpha_{t-1})I\right)$
@@
8:      $\mu_{\theta}(x_{t}):=\sqrt{\frac{1}{\alpha_{t}}}\left(x_{t}-\sqrt{\beta_{t}}\frac{\epsilon_{\theta}(x_{t})}{\sqrt{1-\alpha_{t}}}\right)$
@@ 4.3.3 Multiple Prediction
Diffusion is a noisy process that can generate different results given the same context. In image generation this is a desirable characteristic as the framework can generate different results given the same prompt, increasing the diversity of the generated data. As shown in Figures 4 and 10, SceneSense has the same behavior as these networks and can generate different reasonable predictions based on the same input information. Further there is no compute time increase (assuming enough compute is available) as multiple predictions can be done in parallel. We present our method for merging these various occupancy maps in Section 6.
@@ 5 Robotic System Design
Our robotic system is a quadruped (Spot) as shown in Figure 3 combined with an off-board high performance computer to handle computationally expensive requests. Much of the system design is inherited from the University of Colorado at Boulder MARBLE team that competed in the DARPA Subterranean challenge [19, 7]. For convenience, we provide relevant design decisions in this section and a block diagram of the system in Figure 2.
@@ 5.2 Localization
Localization is required for Spot to perform volumetric mapping. We have implemented the popular LIDAR-based localization method LIO-SAM [47] to provide localization at runtime. To improve localization performance both the IMU and lidar sensor were fastened to a 6061 aluminum sensor plate. This allowed for a high-precision relative transform between the sensors, reducing the necessity of extrinsic calibration. Further LIO-SAM requires aligned sensor timestamps and sensor publish rates to be constant. The LORD IMU however allows for large fluctuation in publish rate due to the USB transmission delays. TO reduce the sensitivity the IMU timestamps are adjusted when messages are not received within $15\%$ of the nominal rate. Additionally, the lidar sensor uses Precision Time Protocol (PTP) to synchronize with the onboard computer. These modifications allowed for consistent localization using LIO-SAM.
@@ 5.3 Frontier Identification and Evaluation
The first work in this domain [13] focused on occupancy prediction around the robot. Later work [14] updates the occupancy prediction framework to predict occupancy at any point in the map, demonstrating enhanced utility for predictions at range. We leverage this enhancement, identifying interesting areas for prediction using the graph-based exploration planner GBPlanner [18]. GBPlanner builds a graph where nodes are potential exploration points and edges are feasible paths to navigate from node to node. A ray casting algorithm is run at each node in the graph to quantify the number of observed, free, and unknown voxels in that node’s field of view. After finding the shortest path to each node the Exploration Gain can be calculated for each node in the graph as follows:
@@ 5.4 Mapping
We use the probabilistic volumetric mapping method OctoMap [17] as the system’s mapping framework. OctoMap was adopted for its log-odds update method for predicting occupied and unoccupied cells. This approach allows for computationally efficient fusion of observed occupancy and predicted occupancy maps. Further discussion on map fusion is provided in Section 6.
000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.6000.10.10.10.10.20.20.20.20.30.30.30.3IoUProbability
000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.6000.10.10.10.10.20.20.20.20.30.30.30.3IoUProbability
@@ 6 Probabilistic Map Merging
In previous work, predicted occupancy is merged into the running occupancy map in a “fire and forget” approach [13]. A given occupancy prediction is temporarily merged into the running occupancy map by setting the predicted occupied cells to 1. Then, when a new occupancy prediction is generated, the previous prediction is removed from the running map and the new prediction is merged in the same way. While this approach is effective in some applications, it limits the ability to accurately maintain a coherent and continuous understanding of the environment. To address these issues, we modify the probabilistic occupancy update formula presented for the OctoMap mapping framework [17].
We define the probability that a voxel $m$ is occupied given an occupancy prediction $d_{t}$ or sensor reading $z_{t}$ as $P(m|d_{t})$ or $P(m|z_{t})$ respectively. The set of sensor estimates $z_{1:t}$ and diffusion estimates $d_{1:t}$ populate the joint set $\{z_{1:t},d_{1:t}\}$ which we denote as $j_{1:t}$. As discussed in [13], SceneSense only operates on voxels $m$ that are not contained in the observed set $\mathcal{O}$, where $z_{t:t-1}=\varnothing$, and therefore $P(m|j_{1:t})$ will never require an update given $P(m|d_{t})$ and $P(m|z_{t})$ at the same time. As such we generate the piece-wise update rule for merging diffusion into the running occupancy map.
@@
$\begin{cases}\left[1+\frac{1-P(m\mid d_{t})}{P(m\mid d_{t})}\frac{1-P(m\mid j_{1:t-1})}{P(m\mid j_{1:t-1})}\frac{P(m)}{1-P(m)}\right]^{-1}&\text{if }m\notin\mathcal{O}\\ \left[1+\frac{1-P(m\mid z_{t})}{P(m\mid z_{t})}\frac{1-P(m\mid j_{1:t-1})}{P(m\mid j_{1:t-1})}\frac{P(m)}{1-P(m)}\right]^{-1}&\text{if }m\in\mathcal{O}.\end{cases}$
In this framework $P(m|z_{t})$ and $P(m|d_{t})$ can be configured to different values prior to runtime. Generally $P(m|d_{t})$ given a predicted occupied cell is set lower than $P(m|z_{t})$ given a sensed occupied cell, as we trust the sensor more than our generative model. By using this probabilistic approach to map merging, the final merged map benefits from prediction persistence as the system explores as well as increased map fidelity due to multi-prediction occupancy refinement.
@@ 7 Occupancy Prediction Experiments and Evaluations
In this section, we provide results and evaluations of the modified SceneSense occupancy prediction framework onboard a real-world system. In particular we evaluate if the SceneSense enhanced occupancy maps better represent a set of fully observed ground-truth maps.
@@
We implement the same denoising network structure presented in Reed et al. [13]. It is a U-net constructed from the HuggingFace Diffusers library of blocks [41] and consists of Resnet [42] downsampling/upsampling blocks. The diffusion model is trained using randomly shuffled pairs of ground truth local occupancy maps $x$. We use Chameleon cloud computing resources [48] to train our model on one A100 with a batch size of $32$ for $250$ epochs or $88,208$ training steps. We use a cosine learning rate scheduler with a 500 step warm up from $10^{-6}$ to $10^{-4}$. The noise scheduler for diffusion is set to $1000$ noise steps.
@@ 7.1 SceneSense Generative Occupancy Evaluation
Figure 8: (a) The robot’s position during exploration near a glass railing, an adversarial scenario for planners that do not consider ground traversability. (b) The running occupancy map shows that the ground suddenly ends to the right of the robot, where the railing prevents it from navigating off the edge. Samplers that do not check for traversability—effective in solving scenarios in Figures 7 and 6—could mistakenly allow paths through open space. Here, SceneSense correctly generates red voxels to indicate untraversable ground, preventing unsafe navigation.
@@
Experimental Setup.   SceneSense predictions are evaluated in 2 environments. Environment 1 is a long hallway with cutouts for classrooms and 1 right turn. Select frames shown in Figures 6 and 7 are from Environment 1. Environment 2 consists of similar carpeted area with 4 hallways and 4 turns, forming a square shape. We evaluate the occupancy prediction framework using the following test configurations.
@@
  3. 3.
    One Shot Map Merging or Probabilistic Map Merging: One shot map merging (OSMM) simply takes the current local occupancy map and a SceneSense occupancy prediction and populates the predicted occupancy information in the running map. Probabilistic map merging (PMM) keeps a running local merged occupancy map that uses update Equation 7 to update the occupancy map for every new occupancy prediction. In practice, each pose will receive 3-5 SceneSense predictions to merge into the running map.
Occupancy Prediction Metrics.   Following similar generative scene synthesis approaches [38, 49] we employ the Fréchet inception distance (FID) [50] and the Kernel inception distance [51] (KID ×$\times 1000$) to evaluate the generated local occupancy grids using the clean-fid library [52]. Generating good metrics to evaluate generative frameworks is a difficult task [53]. FID and KID have become the standard metric for many generative methods due to their ability to score both accuracy of predicted results, and diversity or coverage of the results when compared to a set of ground truth data. While these metrics are fairly new to robotics, which traditionally evaluates occupancy data with metrics like accuracy, precision and IoU, these metrics have been shown to be an effective measure for evaluating predicted scenes [13, 38].
Table 1: Comparison of Running Occupancy (BL) vs. SceneSense (SS). Evaluations of each method are provided as robot-centric (RC) and frontier-centric (FC) generations.
Method Env. 1 Env. 2
FID $\downarrow$ KID×$\times$1000 $\downarrow$ FID $\downarrow$ KID×$\times$1000 $\downarrow$
@@
SS-FC-OSMM 104.2 66.3 133.4 104.4
SS-FC-PMM 30.1 10.3 34.5 9.0
Results Discussion.   The results in Table 1 show that RC predictions are quite similar between OSMM and PMM approaches, reducing the FID of the environments by an average of $28.5\%$ and $25\%$ for OSMM and PMM, respectively. These results are similar to the simulation-based results presented in [13]. However, the FC results show that PMM achieves a much greater improvement than OSMM. This improvement is reflected in the average FID, where a lower value indicates better performance. PMM reduces the average FID by $75\%$, whereas OSMM achieves only an $11\%$ reduction.
Interestingly, The KID values are nearly identical between between SS-RC-PMM and SS-FC-PMM, indicating that the model occupancy predictions at range are as reasonable as the predictions made around the robot, even though there is less information for the predictions at range. KID is known to be less sensitive to outliers when compared to FID [51]. It is likely that the unreasonable predictions that can occur when performing FC predictions are better filtered out by the KID metric, resulting in similar scores.
Figure 10: Example traversal of GBPlanner + Frontier SS. At startup local diffusion is run to allow for planning from startup. Once a graph is constructed and a path is planned the robot begins both traversing the path and making predictions of the frontier occupancy (where predictions are shown in grey). Upon reaching its destination, the graph is extended based on the observed and predicted occupancy and the robot continues exploration.

### v2

@@ 4 SceneSense Design and Implementation
SceneSense [reed2024scenesense, reed2024online] is a diffusion model designed to generate 3D occupancy predictions given an observed occupancy map. Notably, SceneSense only generates occupancy predictions in space that has not been observed, and therefore maintains the fidelity of the observed map. In this section we will discuss the design of the SceneSense network as an unconditional diffusion model as well as our method for training. Then we will discuss our inference time control method of occupancy inpainting.
@@ 4.1.1 Denoising Network
We design our denoising network as an unconditional diffusion model. Specifically the denoising network is a U-net constructed from the huggingface diffusers library of blocks [von-platen-etal-2022-diffusers] and consists of Resnet [he2015resnet] downsampling/upsampling blocks. Notably, as discussed in reed2024scenesense, reed2024online the addition of conditioning for guidance, as often seen in text-to-image diffusion models [rombach2022high], does not generate more accurate predictions and increases inference times in this particular application. These ablations are presented in Appendix A and B.
@@ 4.1.2 Occupancy Mapping
Occupancy mapping allows platforms to build a running map of areas that have been measured to contain matter using onboard sensors like lidar or RGB-D cameras. For our framework we use the popular occupancy mapping framework OctoMap [hornung13octomap] to generate an occupancy map as the platform explores the environment. Importantly, OctoMap provides a probability of occupancy $o\in[0,1]$ for every voxel in the map that has been observed using pose ray casting. This means that as we explore we will maintain not only a map of occupied areas $M_{o}$, but also a map of areas that have been measured to not contain any data $M_{u}$. These maps will later be used to inform SceneSense where occupancy predictions should be made.
@@ 4.2.1 Data Augmentation
In previous SceneSense implementations [reed2024scenesense, reed2024online], the primary area for prediction was centered around the robot. In this paradigm, the SceneSense training data was collected as local occupancy data collected around the robot. However, this training approach struggles to capture the variety of poses possible when doing occupancy predictions at the frontier. The graph-based planner, where poses are selected, can be much closer to structures like walls, and at a wider variety of rotation poses. To account for this, we augment the training data to more closely approximate that of the graph based poses. For each pose we generate 10 noisy pose samples where noise is added from a uniform distribution of [1,1][-1,1] meters in the x,y direction and $0,2\pi]$ in the rotational plane. This modification increases the training set by 10xx but yields much more consistent results at inference time when performing occupancy prediction around frontier nodes in the graph.
@@ 4.2.2 Occupancy Corruption
To corrupt each ground truth local occupancy map $x$ to train the network we add Gaussian noise to $x$ to generate $x_{t}$. This corruption process is defined in 3 where the intensity of the noise is controlled by $\alpha_{t}$ which is configured by a linear noise scheduler [ho2020denoising].
@@ 4.2.3 Loss Function
The network $f_{\theta}$ is trained using the calculated $l_{2}$ loss between the denoised $x_{t}$ prediction and the associated ground truth data $x$. $l_{2}$ loss is a popular diffusion loss function, however other loss functions such as cross-entropy loss or mean squared error can be applied and have had some success in similar diffusion frameworks [glide-nichol22a, rombach2022high, chen2022diffusiondet, ji2023ddp].
@@ 4.3.2 Occupancy Inpainting
Our method of occupancy inpainting is provided in Algorithm 1 which ensures observed space is never overwritten with SceneSense predictions. Inspired by image inpainting methods seen in image diffusion [rombach2022high] and guided image synthesis methods [meng2022sdedit], occupancy inpainting continuously applies the known occupancy information to the diffusion target during inference. To perform occupancy inpainting we select our target region $x$ as a submap of the full occupancy map. From $x$ we build an occupied map $M_{o}$ and an unoccupied map $M_{u}$ composed of observed occupied voxels and observed unoccupied voxels, respectively. From $M_{u}$ and $M_{o}$ we generate noisy representations commensurate with the current diffusion step $t$ using Eq. 3, $M_{un}$ and $M_{on}$, respectively. Finally the data from $M_{un}$ and $M_{on}$ replaces the data in the diffusion target at the associated coordinates.
@@
5:   $x_{obs}^{t-1}\sim\mathcal{N}(\sqrt{\alpha_{t-1}}x_{obs},(1-\alpha_{t-1})I)$
@@
8:   $\mu_{\theta}(x_{t}):=\sqrt{\frac{1}{\alpha_{t}}}\left(x_{t}-\frac{\sqrt{\beta_{t}}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}(x_{t})\right)$
@@ 4.3.3 Multiple Prediction
Diffusion is a noisy process that can generate different results given the same context. In image generation this is a desirable characteristic as the framework can generate different results given the same prompt, increasing the diversity of the generated data. As shown in Figures 4 and 9, SceneSense has the same behavior as these networks and can generate different reasonable predictions based on the same input information. Further there is no compute time increase (assuming enough compute is available) as multiple predictions can be done in parallel. We present our method for merging these various occupancy maps in Section 6.
@@ 5 Robotic System Design
Our robotic system is a quadruped (Spot) as shown in Figure 3 combined with an off-board high performance computer to handle computationally expensive requests. Much of the system design is inherited from the University of Colorado at Boulder MARBLE team that competed in the DARPA Subterranean challenge [biggie2023flexible, chung2023into]. For convenience, we provide relevant design decisions in this section and a block diagram of the system in Figure 2.
@@ 5.2 Localization
Localization is required for Spot to perform volumetric mapping. We have implemented the popular LIDAR-based localization method LIO-SAM [shan2020liosam] to provide localization at runtime. To improve localization performance both the IMU and lidar sensor were fastened to a 6061 aluminum sensor plate. This allowed for a high-precision relative transform between the sensors, reducing the necessity of extrinsic calibration. Furthermore, LIO-SAM requires aligned sensor timestamps and sensor publish rates to be constant. The LORD IMU however allows for large fluctuation in publish rate due to the USB transmission delays. To reduce the sensitivity the IMU timestamps are adjusted when messages are not received within $15\%$ of the nominal rate. Additionally, the lidar sensor uses Precision Time Protocol (PTP) to synchronize with the onboard computer. These modifications allowed for consistent localization using LIO-SAM.
@@ 5.3 Frontier Identification and Evaluation
The first work in this domain [reed2024scenesense] focused on occupancy prediction around the robot. Later work [reed2024online] updates the occupancy prediction framework to predict occupancy at any point in the map, demonstrating enhanced utility for predictions at range. We leverage this enhancement, identifying interesting areas for prediction using the graph-based exploration planner GBPlanner [dang2020graph]. GBPlanner builds a graph where nodes are potential exploration points and edges are feasible paths to navigate from node to node. A ray casting algorithm is run at each node in the graph to quantify the number of observed, free, and unknown voxels in that node’s field of view. After finding the shortest path to each node the Exploration Gain can be calculated for each node in the graph as follows:
@@ 5.4 Mapping
We use the probabilistic volumetric mapping method OctoMap [hornung13octomap] as the system’s mapping framework. OctoMap was adopted for its log-odds update method for predicting occupied and unoccupied cells. This approach allows for computationally efficient fusion of observed occupancy and predicted occupancy maps. Further discussion on map fusion is provided in Section 6.
@@ 6 Probabilistic Map Merging
In previous work, predicted occupancy was merged into the running map in a “fire and forget” fashion [reed2024scenesense]. Each prediction was incorporated by setting its occupied cells to 1, but when a new prediction was generated, the previous one was discarded and replaced. While simple, this approach prevents the system from maintaining a coherent and continuous estimate of the environment. A single out of distribution prediction can temporarily warp the map, causing the planner to divert the robot along a spurious path. Once the next prediction arrives and overwrites the first, the planner abruptly redirects, potentially producing oscillatory shifts in the planned trajectory. To address these issues, we modify the probabilistic occupancy update formula presented for the OctoMap mapping framework [hornung13octomap].
We define the probability that a voxel $m$ is occupied given an occupancy prediction $d_{t}$ or sensor reading $z_{t}$ as $P(m|d_{t})$ or $P(m|z_{t})$, respectively. The set of sensor estimates $z_{1:t}$ and diffusion estimates $d_{1:t}$ populate the joint set $\{z_{1:t},d_{1:t}\}$ which we denote as $j_{1:t}$. As discussed in [reed2024scenesense], SceneSense only operates on voxels $m$ that are not contained in the observed set $\mathcal{O}$, where $z_{t:t-1}=\varnothing$, and therefore $P(m|j_{1:t})$ will never require an update given $P(m|d_{t})$ and $P(m|z_{t})$ at the same time. As such we generate the piece-wise update rule for merging diffusion into the running occupancy map.
@@
[…]
In this framework $P(m|z_{t})$ and $P(m|d_{t})$ can be configured to different values prior to runtime. Generally, $P(m|d_{t})$ given a predicted occupied cell is set lower than $P(m|z_{t})$ given a sensed occupied cell, as we trust the sensor more than our generative model. By using this probabilistic approach to map merging, the final merged map benefits from prediction persistence as the system explores as well as increased map fidelity due to multi-prediction occupancy refinement.
@@ 7 Occupancy Prediction Experiments and Evaluations
In this section, we provide results and evaluations of the modified SceneSense occupancy prediction framework onboard a real-world system. In particular we evaluate if the SceneSense enhanced occupancy maps better represent a set of fully observed ground truth maps.
@@
We implement the same denoising network structure presented in reed2024scenesense. It is a U-net constructed from the HuggingFace Diffusers library of blocks [von-platen-etal-2022-diffusers] and consists of Resnet [he2015resnet] downsampling/upsampling blocks. The diffusion model is trained using randomly shuffled pairs of ground truth local occupancy maps $x$. We use Chameleon cloud computing resources [keahey2020chameleon] to train our model on one A100 with a batch size of $32$ for $250$ epochs or $88,208$ training steps. We use a cosine learning rate scheduler with a 500 step warm up from $10^{-6}$ to $10^{-4}$. The noise scheduler for diffusion is set to $1000$ noise steps.
@@
Evaluation Metrics. We evaluate predicted occupancy using a classical computer vision metric, Intersection over Union (IoU) [everingham2010pascal], and newer generative modeling metrics, Fréchet Inception Distance (FID) [heusel2017gansFID] and Kernel Inception Distance (KID) [binkowski2018demystifyingKID]. IoU provides a direct measure of overlap between predicted and ground truth occupied voxels. To assess the realism and diversity of generated occupancy predictions, we report FID and KID. FID measures the Fréchet distance between Gaussian statistics of features extracted from predicted and real samples, while KID uses a polynomial kernel Maximum Mean Discrepancy, rather than assuming Gaussian distributions, to provide an unbiased estimate. Both metrics have become standard in generative modeling, and we compute them using the clean-fid library [parmar2021cleanfid].
It is worth noting that generating good metrics to evaluate generative frameworks remains a difficult task [naeem2020reliable]. FID and KID have become standard metrics for many generative methods due to their ability to score accuracy of predicted results and diversity or coverage of the results when compared to a set of ground truth data. While these metrics are fairly new to robotics, which traditionally evaluates occupancy data with metrics like accuracy, precision and IoU, we report FID and KID, in addition to IoU, because they have been shown to be an effective measure for evaluating predicted scenes [reed2024scenesense, tang2023diffuscene].
@@ 7.1 SceneSense Generative Occupancy Evaluation
Figure 7: (a) The robot’s position during exploration near a glass railing, an adversarial scenario for planners that do not consider ground traversability. (b) The running occupancy map shows that the ground suddenly ends to the right of the robot, where the railing prevents it from navigating off the edge. Samplers that do not check for traversability—effective in solving scenarios in Figures 6 and 5—could mistakenly allow paths through open space. Here, SceneSense correctly generates red voxels to indicate untraversable ground, preventing unsafe navigation.
@@
Experimental Setup. SceneSense predictions are evaluated in 2 environments. Environment 1 is a long hallway with cutouts for classrooms and 1 right turn. Select frames shown in Figures 5 and 6 are from Environment 1. Environment 2 consists of similar carpeted area with 4 hallways and 4 turns, forming a square shape. We evaluate the occupancy prediction framework using the following test configurations.
@@
    Robot-centric or Frontier-centric: Robot-centric diffusion (RC) predicts occupancy at a radius of $3.3m$ about the robot while frontier-centric diffusion (FC) predicts occupancy at a radius of $3.3m$ at an identified location in the map, which has a maximum range of 7mm from the robot.
  3. 3.
    One Shot Map Merging or Probabilistic Map Merging: One shot map merging (OSMM) simply takes the current local occupancy map and a SceneSense occupancy prediction and populates the predicted occupancy information in the running map using the “fire and forget” approach described in Section 6. Probabilistic map merging (PMM) keeps a running local merged occupancy map that uses update Equation 7 to update the occupancy map for every new occupancy prediction. In practice, each pose will receive 3-5 SceneSense predictions to merge into the running map.
Generative Occupancy Results Discussion. The results in Table 1 show that RC predictions are quite similar between OSMM and PMM approaches, reducing the FID of the environments by an average of $28.5\%$ and $25\%$ for OSMM and PMM, respectively. These results are similar to the simulation-based results presented in [reed2024scenesense]. However, the FC results show that PMM achieves a much greater improvement than OSMM. This improvement is reflected in the average FID, where a lower value indicates better performance. PMM reduces the average FID by $75\%$, whereas OSMM achieves only an $11\%$ reduction.
Table 1: Comparison of Running Occupancy (BL) vs. SceneSense ($S$). Evaluations of each method are provided as robot-centric (RC) and frontier-centric (FC) generations.
Method Env. 1 Env. 2
@@
$S$-FC-OSMM 104.2 66.3 133.4 104.4
$S$-FC-PMM 30.1 10.3 34.5 9.0
Interestingly, The KID values are nearly identical between SS-RC-PMM and SS-FC-PMM, indicating that the model occupancy predictions at range are as reasonable as the predictions made around the robot, even though there is less information for the predictions at range. KID is known to be less sensitive to outliers when compared to FID [bińkowski2018demystifying]. The unreasonable predictions that can occur when performing FC predictions are likely better filtered out by the KID metric, resulting in similar scores.
Figure 9: Example traversal of GBPlanner + Frontier $S$. At startup local diffusion is run to allow for immediate planning. Once a graph is constructed and a path is planned the robot begins both traversing the path and making predictions of the frontier occupancy (where predictions are shown in grey). Upon reaching its destination, the graph is extended based on the observed and predicted occupancy and the robot can continue exploring.
@@
We see an expected increase in FID and KID for frontier-centric prediction. Around the robot, more of the space has been directly observed with lidar sensor measurements. This means less unknown voxels are available for generative prediction since SceneSense does not override occupancy values for voxels that are observed occupied or free based on sensor measurements. As a result, the underlying structure of the environment is more similar between predictions and the ground truth. This results in lower FID and KID values for robot-centric. When making predictions at the frontier, more unknown voxels are available, leading to larger variance in the predictions. This can result in a larger difference between the ground truth map and the predicted occupancy generated with SceneSense, leading to larger FID and KID values.